OpenStudy (nubeer):

Calc 3... how do we find the integration factor of equation which is not exact.. (i want to know for in which both x and y is used as integration factor)

OpenStudy (unklerhaukus):

do you have an example we can work through /

OpenStudy (nubeer):

hmm no i don't have any example .. we were asked to look for it.. i just want to know any general method.. any link or inofrmation would appreciate it.. its basically part of diffrential equations.

OpenStudy (unklerhaukus):

\[\left(3x+\frac 6y\right)\text dx+\left(\frac {x^2}{y}+\frac{3y}{x}\right)\text dy =0\]

OpenStudy (unklerhaukus):

\[\frac{\partial M}{\partial y}=\frac {-6}{y^{2}}\qquad\qquad \frac{\partial N}{\partial x}=\frac{2x}{y}-\frac{3y}{x^2}\]\[\qquad\qquad\frac{\partial M}{\partial y}\neq \frac{\partial N}{\partial x}\]

OpenStudy (nubeer):

ok..

OpenStudy (unklerhaukus):

\[R=R(x,y)=x^my^n\]

OpenStudy (nubeer):

ok this is soemthing new.. is this the general form of the solution?

OpenStudy (unklerhaukus):

im using R for the integrating factor

OpenStudy (nubeer):

ok..

OpenStudy (unklerhaukus):

\[\frac{\partial R(x,y)M}{\partial y}\quad= \frac{\partial R(x,y)N}{\partial x}\] using the product rule \[R_y(x,y)M+R(x,y)\frac{\partial M}{\partial y}=R_x(x,y)N+R(x,y)\frac{\partial N}{\partial x}\]

OpenStudy (nubeer):

ok i get this much

OpenStudy (unklerhaukus):

\[\small R_y(x,y)\left(3x+\frac 6y\right)+R(x,y)\left(\frac {-6}{y^{2}}\right)=R_x(x,y)\left(\frac {x^2}{y}+\frac{3y}{x}\right)+R(x,y)\left(\frac{2x}{y}-\frac{3y}{x^2}\right)\] \[\small nx^my^{n-1}\left(3x+\frac 6y\right)+x^my^n\left(\frac {-6}{y^{2}}-\frac{2x}{y}+\frac{3y}{x^2}\right)-mx^{m-1}y^n\left(\frac {x^2}{y}+\frac{3y}{x}\right)=0\]

OpenStudy (unklerhaukus):

.... somehow find n and m ,

OpenStudy (nubeer):

how? any particular method?

OpenStudy (unklerhaukus):

something to do with the matching indices

OpenStudy (nubeer):

hmm still dont get the last part..

OpenStudy (unklerhaukus):

me either

OpenStudy (nubeer):

hmm ok its fine but thanks for the help bro.. :) will look on internet

OpenStudy (unklerhaukus):

if you find a really good sight can you tell me please

OpenStudy (nubeer):

sure i will let u know..

OpenStudy (nubeer):

http://www.cliffsnotes.com/study_guide/Integrating-Factors.topicArticleId-19736,articleId-19711.html @UnkleRhaukus this is a good one but it dont tell for the both x and y .. will post it if i find it